OpenStudy (anonymous):

A collection of 27 coins, all nickels and dimes, is worth $2.10. How many of each coin are there?

OpenStudy (anonymous):

5n+10d=210

OpenStudy (anonymous):

n=number of nickels, d=number of dimes

OpenStudy (anonymous):

no because there is a total of 27 coins, not 15.

OpenStudy (anonymous):

missed that part .. hang on :)

OpenStudy (anonymous):

We need to solve this : n+d=27 5n+10d=210

OpenStudy (anonymous):

d=27-n 5n+10(27-n)=210 5n+270-10n=210 -5n=-60 n=12 d=27-n=27-12=15 so there are 12 nickels and 15 dimes

OpenStudy (anonymous):

and here are the 27 coins :)

OpenStudy (anonymous):

yayy! :) thank you so much...Btw i was trying to figure it out too.

OpenStudy (anonymous):

np :)

OpenStudy (amistre64):

n = (2.10 - 27(.10))/(.05-.10) d = (2.10 - 27(.05))/(.10-.05)

OpenStudy (amistre64):

so yeah, 12 ns and the rest are ds :)